OpenStudy (anonymous):

solve this system of equations x+y+z=-2 2x+5y+2z=2 -x+7y-3z=24

hero (hero):

Add the first and third equations together to make it easier

hero (hero):

Than add that result to the middle equation

hero (hero):

Then solve for y

OpenStudy (radar):

y=2, x=-1, z=-3 Use elimination method.

OpenStudy (radar):

Many ways to approach, but you can solve for y by taking the second equation and subtracting the first equation multiplied by 2.

hero (hero):

Yeah, my approach didn't help much

OpenStudy (radar):

2x +5y + 2z = 2 2x +2y + 2z = -4 ----------------- 3y = 6 y=2 Now you take it from there.

OpenStudy (radar):

Hero it did eliminate one of the unknowns, so it was a good start.

OpenStudy (radar):

Now substitute for y from what you got with Hero suggestion. 8y-2z=22 16-2z=22 -2z=6 z=-3 Now you are home free.